OpenStudy (anonymous):

I need some help getting started on this problem. Find the cube roots of 9√3 + 6i.

OpenStudy (anonymous):

I start with De Moivre's Theorem? z = r(cosθ + i sinθ)

OpenStudy (anonymous):

r^2 = (9√3)^2 + (6)^2 r^2 = (243 + 36) r = 16.70329309... seems messy, did I screw up?

OpenStudy (anonymous):

I want to make sure I'm going about this correctly so far before proceeding.